OpenStudy (anonymous):

What pressure, in millimeters of mercury (mm Hg), is equivalent to 2.13 atmospheres?

OpenStudy (anonymous):

1 atmosphere = 760.0 mmHg

OpenStudy (abb0t):

@yummydum is correct. Use dimensional analysis to cross cancel units, just as you did in your algebra class with fractions. In this case, your variables are units. For instance: \[\frac{ a }{ 2b } \times \frac{ 4b }{ a } = 2b\]

OpenStudy (anonymous):

now just set up a proportion:\[{{1atm}\over{760.0mmHg}}={{2.13atm}\over{\color{red}{x}~mmHg}}\]and solve for x

OpenStudy (anonymous):

right but they have set the choices up like: 3.57 x 10^ 2 mm Hg

OpenStudy (anonymous):

3.57 x 10^2 mm Hg

OpenStudy (anonymous):

that just means 357 mmHg